OpenStudy (anonymous):

please help me this is impotent i have this assessment that is 15 questions long I'm on the 8th one Jake needs 9 pieces of rope that are each the same length. He cuts the pieces from a rope that is 8 yards long and uses up the whole piece of rope. How long is each piece of rope?

OpenStudy (anonymous):

fractions are so hard please help this is super hard I'm only in 4 grade using my moms email

OpenStudy (anonymous):

will any one please help me i will become fan i only have two fans because i begun open study today help me please

OpenStudy (yoloshroom):

does the answer need to be in feet? or yards?

OpenStudy (anonymous):

yards

OpenStudy (yoloshroom):

so roughly .89 yards per rope. we find this by dividing 8 by 9. so 8/9 = .8999999 so we can say 8/9 = .89 roughly. so .89yd per rope

OpenStudy (futureboy1):

what measure are u using?

OpenStudy (anonymous):

but i need it in fractions do you think you can do it in fractions?

OpenStudy (anonymous):

im working with fractions in math

OpenStudy (futureboy1):

@YoloShroom placed a fraction

OpenStudy (anonymous):

.so its 8/9 the answer is 8 over 9???

OpenStudy (futureboy1):

wait don move on yet

OpenStudy (anonymous):

im not

OpenStudy (yoloshroom):

Is your question able to be put in electronically? if so, it's 8/9 because easch piece is .888888888889 or .89yd each. if that doesnt work swap it to 9/8 because math is op.

OpenStudy (anonymous):

math is so hard you know i think the answer is 8/9

OpenStudy (futureboy1):

hold on almost got the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for helping

OpenStudy (futureboy1):

i think it would be 2 7/10 I think: 8 yrd= 24 ft 24 ft = 288 in 288 divided by 9=32 in 32 in = 2.7 ft.....and 2.7 ft as a fraction = 2 7/10

OpenStudy (futureboy1):

Thats my guess at it but it's your choice?

OpenStudy (futureboy1):

@crianca

OpenStudy (anonymous):

thanks

OpenStudy (futureboy1):

but on my test why did u say B was the answer

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i do believe so

OpenStudy (futureboy1):

can u show me the math?

OpenStudy (anonymous):

i need help with another problem Express the quotient in simplest form. 3 divided by 1/4 =m i need it done quikly i have been working on this math assessment for an hour now

OpenStudy (futureboy1):

give me a medal

OpenStudy (anonymous):

i have troble with math i tryed to figure it out but had a hard time i dont know how to give medels

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

please will you help me

OpenStudy (futureboy1):

click best response

OpenStudy (anonymous):

ok thanks there you go

OpenStudy (anonymous):

do you think i should change my picture? i got lots of pictures of snakes, catie,and dragons